OpenStudy (anonymous):

A line has a slope of 0.5 and a y-intercept of (0, -2). What is the value of y when x = 4?

OpenStudy (anonymous):

slope intercept gotta be the easiest way to get the equation of the line ^.^ the equation is y = 0.5x - 2

OpenStudy (anonymous):

Thank you ~!

OpenStudy (anonymous):

first deduce equation of line y=mx+c where m is slope c is intercept y=0.5x-2 put x=4 y=2-2 y=0